A poll found that 74​% of a random sample of 1013 adults said that they believe in ghosts. a. Determine the margin of error for a​ 99% confidence interval. (Round to three decimal places) E=
Answer #1

Solution :  

Given that,

n = 1013

Point estimate = sample proportion = = 74% = 0.74

1 - = 1 - 0.74 = 0.26

At 99% confidence level the z is ,

= 1 - 99% = 1 - 0.99 = 0.01

/ 2 = 0.01 / 2 = 0.005

Z/2 = Z0.005 = 2.576

Margin of error = E = Z / 2 * (( * (1 - )) / n)

= 2.576 * (((0.74 * 0.26) / 1013)

E = 0.036